The stones tied on top are evenly rotating in a vertical plane. Upholstery
the difference between the maximum and minimum values of the pulling force of the wire
If ∆T = 10N
Find the mass of the stone m.
1
Expert's answer
2020-03-25T11:07:35-0400

At the top:


"T_t=m\frac{v_t^2}{r}-mg,T_b=m\frac{v_b^2}{r}+mg"

From the conservation of energy:


"0.5mv_t^2+2mgr=0.5mv_b^2"

"T_b-T_t=6mg"

"10=6(9.8)m"

"m=0.17\ kg"